OpenStudy (anonymous):

I need help with my fractions homework. It is 5th grade math.

OpenStudy (sandra):

hey =). ok so give an example problem to give me an idea of what you're struggling with

OpenStudy (anonymous):

Well, I am Comparing and Ordering fractions, Mixed Numbers, and Decimals.

OpenStudy (sandra):

ok sure

OpenStudy (anonymous):

5 1/3 and 6.7

OpenStudy (sandra):

right, ok

OpenStudy (anonymous):

Can you explain it to me.

OpenStudy (anonymous):

hahahah homeworkhelpgirl... made a new account

OpenStudy (sandra):

so 5 1/3 means 5 , and then 1/3 of 1

OpenStudy (sandra):

in decimal form, that means 5.333333333

OpenStudy (anonymous):

Remember I am only in the 5th grade,,,,

OpenStudy (anonymous):

sandra can you help me with regression when you're finished please?

OpenStudy (sandra):

but conceptually, you can see that 5 1/3 is less than six, because 5 1/3 = 5 + 1/3 (which is less than one), and 6.7 means 5 + 1.7 (which is more than one)

OpenStudy (sandra):

sorry briibaby, too advanced for me!

OpenStudy (anonymous):

What?

OpenStudy (sandra):

another way to compare them is to turn them both into fractions

OpenStudy (sandra):

but let's see some other sample problems too

OpenStudy (sandra):

for example, what is bigger, 1/2 or 1/3?

OpenStudy (anonymous):

1/3?

OpenStudy (sandra):

nope, it's just the bottom number is bigger. let me show you how to compare

OpenStudy (sandra):

to compare fractions, it helps to find what's called a common denominator

OpenStudy (sandra):

for example, if I asked you what's bigger, 1/3, or 2/3, what would you say?

OpenStudy (anonymous):

Wait :) I have some numbers with decimals with them!

OpenStudy (sandra):

ok but 2/3 is bigger than 1/3, because the top number is bigger in 2/3 than 1/3 - and they're fractions out of the same overall unit (3's)

OpenStudy (sandra):

but go ahead and post another example with decimals

OpenStudy (anonymous):

How do convert fractions into decimals?

OpenStudy (sandra):

well, it's easiest to make it so some power of 10 is on the bottom

OpenStudy (sandra):

for example, 1/10 = .1 , 1/100 = .01, 1/1000 = .001 The rule is, you're allowed to move the decimal one to the left with each zero that's on the bottom

OpenStudy (anonymous):

I don't understand.

OpenStudy (sandra):

well, 1/10 is another way of writing "one divided by ten"

OpenStudy (sandra):

and when I divide by some power of 10 - for example 10, 100, 1000, 2000, all that means is moving the decimal point one to the left

OpenStudy (sandra):

so 1 divided by 10 is .1

OpenStudy (sandra):

2 divided by 10 is .2

OpenStudy (sandra):

I just move the point one to the left

OpenStudy (sandra):

multiplying by some power of 10 means I get to move the decimal point one to the right, and add a zero

OpenStudy (sandra):

so 1 x 10 = 10 (added one zero), 2 x 100 = 200 (I get to add two zeros), and 6.7 x 10 = 67 (I get to move the decimal point to the right)

OpenStudy (sandra):

so there are two ways to compare a decimal and a fraction

OpenStudy (sandra):

one way would be to convert the fraction to a decimal - and then compare them

OpenStudy (sandra):

another way would be to convert the decimal to a fraction, and then compare those fractions

OpenStudy (sandra):

I imagine they want you to learn the second approach

OpenStudy (sandra):

now to compare fractions, as I was saying earlier, you need to find what's called a "common denominator"